# Collision-Free Robot Navigation in Crowded Environments using Learning based Convex Model Predictive Control

### 03 Mar 2024 v1

@@ III-E Observation Space Formulation
At time t, the observation for the agent, $o_{t}$, as illustrated in Equation 6 and Figure 5, includes: the Euclidean distance $d_{t}$ from robot to navigation goal, the velocity magnitude $v_{t}$, the angular deviation $d\theta_{t}$ between robot’s velocity direction and the line to goal, the short-term and long-term reference points $Q_{t-1}^{s}$ and $Q_{t-1}^{l}$ from the previous frame’s action, the MPC-optimized points $Q_{1}^{*}$ and $Q_{N}^{*}$ (defined in Section III-D), and the convex region $convex_{t}$.
@@
Due to the variability in shape of the convex static obstacle-free region ($convex_{t}$) derived from different point cloud data, the number of vertices in convex polygons is not constant. This poses a challenge for neural networks, which require input of fixed dimension. To address this issue, the vertex count of the computed convex region, $num_{v}$ , is adjusted to match the predefined number $rnum_{v}$: vertices are added through interpolation if the count $num_{v}$ is less than $rnum_{v}$, or reduced via sparsification if $num_{v}$ exceeds $rnum_{v}$.
The agent can extract motion information about dynamic obstacles and its own historical trajectories from continuous frames implicitly. We define the agent’s state as $s_{t}=\left(o_{t-2},o_{t-1},o_{t}\right)$, a concatenation of observations from three consecutive frames. This state representation, with its dimensionality set to $336rnum_{v}+$ 33, optimizes the processing of temporal and spatial information, enhancing the decision-making process.
@@ III-F Reward Function Formulation
The reward function plays a crucial role by setting the missions for agent. Agent’s strategy is formulated based on the anticipation of future rewards, making reward function a critical mechanism for conveying tasks to the agent.
A positive reward, $r_{success}$, is awarded to the agent when its Euclidean distance to the navigation goal, $d_{t}$, falls below specified threshold, $d_{th}$, indicating successful arrival at the destination.
@@
Collision penalty $r_{t}^{o}$, derived from distance data $scan$ collected by LiDAR, is dynamically adjusted. The penalty decrease exponentially as the distance increases, controlled by a positive decay factor $w_{obs}$.
@@
To mitigate sparse reward, we incorporate potential-based reward shaping. A global path is computed at each episode using the Hybrid A* algorithm [16], based on robot’s starting position, destination, kinematics, and global map. This path, used solely for reward calculation, informs the establishment of the Frenet [17] coordinate for tracking the robot’s position $(\varsigma_{t},l_{t})$, where $\varsigma_{t}$ is the longitudinal displacement and the lateral deviation $l_{t}$. The reward item $r_{t}^{a}$ guides the agent towards the global path direction, while aiming to minimize lateral deviation $l_{t}$, a penalty $r_{t}^{l}$ is introduced, as depicted in Equation 9 and Fig. 6.
$r_{t}^{a}$ =$=\begin{cases}0&\text{if }t=1\\ w_{approach}(\varsigma_{t}-\varsigma_{t-1})&\text{otherwise}\end{cases}$ (9)
$r_{t}^{l}$ $=w_{close}{l_{t}}^{2}$
Figure 6: Schematic of Global Path Based Potential Shaping Reward
In environment dense with dynamic obstacles, the optimal path diverges significantly from that in static settings, rendering potential shaping reward based on global path ineffective for accurate behavior assessment. To address this, we introduce a potential shaping reward $r_{t}^{a2}$ , based on the Euclidean distance $d_{t}$ between the robot and its destination.
@@
Short-term and long-term reference points, $Q_{t}^{s}$ and $Q_{t}^{l}$, discussed in Section III-C, might not be practical to achieve. However, the MPC-optimized positions $Q_{i}^{*},i=\left\{1,...,N\right\}$, outlined in Section III-D, comply with kinematic constraints and provide a more feasible reference. Thus, to penalize the error between reference points and MPC-optimized points, we introduce $r_{t}^{f}$, which encourages the generation of more reachable reference points.
@@
Significant changes in reference points across frames can lead to discrepancies in both the length and direction of the MPC-optimized trajectory, demanding increased control efforts to counteract the impact of preceding frame. To address this, a penalty, $r_{t}^{c}$, on changes in reference points is introduced. This penalty can also facilitates rapid learning of smooth direction change during the initial training phase.
@@
The final reward function, denoted as $r_{t}$, is composed of above eight components. These components can be adjusted or selectively ignored during phases of training.
$r_{t}=r_{t}^{e}+r_{t}^{s}+r_{t}^{a}+r_{t}^{a2}+r_{t}^{l}+r_{t}^{o}+r_{t}^{c}+r_{t}^{f}$ (13)
@@ III-G Network Architecture
The convex static obstacle-free region efficiently removes redundant information from the raw LiDAR point cloud, including noise, duplicate points, and distant obstacles. This preprocessing step enables the adoption of a more streamlined network architecture for fitting policy and value functions, as depicted in Figure 7. Given that separate networks can yield better performance in practice [18], the policy and value network are updated independently, without sharing parameters.
Figure 7: Network Architecture: The value network $V_{\phi}$ and policy network $\pi_{\theta}$ are structured as four-layer fully connected networks. The policy network is augmented with a log standard deviation parameter ($\ln\sigma_{t}$) for each action dimension. This configuration primarily facilitates the generation of stochastic policies by allowing actions to be sampled from a Gaussian distribution, where $\sigma_{t}$, derived from $\ln\sigma_{t}$, modulates the balance between exploration and exploitation.
@@ IV EXPERIMENTS
In this section, we introduce our training process along with the definition of seven Stages that vary in complexity (Table I). These stages are specifically designed for curriculum learning and subsequent evaluation. Following this, we evaluate our method in three parts:
(1) To demonstrate the advantages of our method in terms of action and state space design, we conduct an ablation study. This study compares our approach with four different DRL navigation methods constructed with various action and state spaces (Section IV-B).
(2) To validate the effectiveness of our method in reward function design, we conduct another ablation study. This study compares our method against three other DRL navigation methods with different reward configurations (Section IV-C).
(3) To further analyze our method’s advantages over non-DRL navigation methods in dynamic and crowded environments, we compare our method with the Timed Elastic Band (TEB) method in test scenarios (Section IV-D).
@@ IV-A Training Procedure
Our method employs a multi-stage training strategy that progresses from simple to complex scenarios, rooted in the principles of curriculum learning, aiming to enhance the agent’s learning efficiency by optimizing the sequence of experience acquisition. Initially, the agent learns basic navigation movements in an obstacle-free environment (Stage 1), and then it gradually encounters more static and dynamic obstacles.
@@ IV-A1 Stage Definition
Table I displays the parameters of static and dynamic obstacles across seven Stages for staged training and subsequent evaluation. Static obstacles are randomly generated polygons with three or four sides, where the maximum area of these polygons does not exceed 2. The dynamic obstacles are circular in shape.
To assess the performance of the trained agent, we generate 1,000 test scenarios for each Stage, from 2 to 7, using random numbers not present in the training dataset to configure the simulator.
@@ IV-A2 Implementation Details
Our work’s leverages the Arena-Rosnav framework [19], developed within the ROS architecture, facilitating robust robotic navigation solutions. We also adopt the implementation of Proximal Policy Optimization (PPO) [20] from Elegant-RL [21], which offers tools for network customization and training process monitoring. The simulation experiments are conducted on a high-performance PC equipped with an Intel Core i7-7700K CPU, an Nvidia RTX 2080Ti GPU, and 32GB of RAM.
@@ IV-B Ablation Study on State and Action Space
To verify the effectiveness of the proposed method in terms of action and state space design, we conducted comparative experiments against four methods:
Design 1: Using the conventional end-to-end architecture where the observation $o_{t}$ includes LiDAR data $scan_{t}^{{}^{\prime}}$ Action is defined as a set of two-dimensional velocities $\left(v_{x},v_{y}\right)$ , with the policy network’s output layer using $tanh$ activation. The mapping from the policy network output $\left(\alpha_{t},\beta_{t}\right)$ to $\left(v_{x},v_{y}\right)$ is detailed in (14).
@@
Design 2: The state observation $o_{t}$ updates to (15), incorporating the convex static obstacle-free region. The action, consisting of $(v_{x},v_{y})$, remains consistent with that of Design 1.
@@
Design 3: The action is simplified to $(Q_{t}^{l})$, focusing exclusively on long-term reference point. The calculation method of a single $(Q_{t}^{l})$ is similar to that of short-term reference point $(Q_{t}^{s})$ in Section III-C, except that single $(Q_{t}^{l})$ is sampled from the long-term action space. Additionally, Design 3 also streamlines the state observation as in (16) by removing terms related to $Q_{t}^{s}$.
@@
Design 4: The observation $o_{t}$ incorporates raw LiDAR data $scan_{t}^{{}^{\prime}}$, as in (17). The action is still $\left(Q_{t}^{l},Q_{t}^{s}\right)$, but calculated within the intersection of the point cloud’s coverage and the robot’s kinematic limits. Given that $scan_{t}^{{}^{\prime}}$ is non-convex and cannot be used as a constraint for convex optimization, the constraint that the optimized trajectory must be within the convex region is removed during the MPC calculation.
@@
After training from Stage 2 to 4, the performance of our method and Designs 1 to 4 in test scenarios is presented in Table II.
@@
Stage Method Success Rate Time (s) Distance (m) Speed (m/s)
Design 1 76% 4.0 11.28 2.87
Design 2 76% 4.0 11.11 2.8
2 Design 3 90.3% 9.0 15.26 1.72
Design 4 83% 5.0 11.66 2.23
Ours 89.2% 5.5 12.19 2.18
Design 1 79.9% 4.0 11.37 2.9
Design 2 79% 4.0 11.4 2.88
3 Design 3 87.8% 8.8 15.42 1.77
Design 4 83.5% 4.9 11.59 2.24
Ours 89.1% 5.9 13.04 2.17
Design 1 80.9% 3.9 11.25 2.95
Design 2 75.1% 3.8 10.78 2.89
4 Design 3 84.5% 9.0 15.13 1.69
Design 4 83.7% 4.8 11.19 2.2
Ours 85.5% 5.8 12.51 2.1
In terms of action space design, Design 3 simplifies action to a single long-term reference point. Although Design 3 has an average navigation success rate of 87.53%, close to our method’s 87.93%, Design 3’s average navigation time is 3.2s longer and the average navigation distance is 2.69m longer than our method in the three Stages. On the other hand, Designs 1 and 2, which utilize continuous two-dimensional velocity for the action space, exhibit success rates lower by 9% and 11.23%, respectively, compared to our method.
In terms of state space design, compared with Design 1 and 4 that directly use raw LiDAR data as the observation, our method’s average navigation success rate is 9% and 4.53% higher, respectively. The ablation study demonstrates the effectiveness of our method in the design of the action and state space.
@@ IV-C Ablation Study on Reward Function
Reward function, the configuration used in our method, incorporates all items.
$r_{t1}=r_{t}^{e}+r_{t}^{s}+r_{t}^{a}+r_{t}^{a2}+r_{t}^{l}+r_{t}^{o}+r_{t}^{c}+r_{t}^{f}$ (18)
@@
Stage Reward Success Rate Time (s) Distance (m) Speed (m/s) Total Abs Acc
5 $r_{t1}$ 87.1% 2.9 4.59 1.57 26475.46
$r_{t2}$ 87.8% 2.9 4.65 1.57 27600.56
$r_{t3}$ 84.8% 3.4 4.38 1.40 27781.34
$r_{t4}$ 86.7% 2.9 4.65 1.58 26460.66
6 $r_{t1}$ 78.1% 2.9 4.12 1.42 19507.60
$r_{t2}$ 77.3% 2.8 4.10 1.43 20397.66
$r_{t3}$ 73.1% 3.4 3.85 1.25 21988.81
$r_{t4}$ 74.6% 3.0 4.35 1.46 19950.77
7 $r_{t1}$ 68.5% 3.1 4.22 1.36 6348.57
$r_{t2}$ 69.9% 3.0 4.15 1.37 7218.53
$r_{t3}$ 67.1% 3.6 3.98 1.19 6080.69
$r_{t4}$ 65.3% 2.9 3.88 1.36 5668.30
In assessing how composite reward function affect the smoothness of motion, Total Absolute Acceleration (Total Abs Acc) serves as a critical metric. Total Abs Acc is the sum of agent’s absolute acceleration ($a_{i}=\frac{|v_{i+1}-v_{i}|}{\Delta t}$, where $\Delta t$=0.1s) taken across steps within successful episodes. A lower Total Abs Acc implies a smoother motion strategy, which is vital for safe and smooth navigation.
The ablation study clarifies the significance of the MPC-optimized points error penalty ($r_{t}^{f}$) for success rates, as seen in $r_{t1}$’s performance, with success rates of 87.1%, 78.1%, and 68.5% across stages. Exclusion of $r_{t}^{f}$ in $r_{t3}$ leads to decreased success rates (84.8%, 73.1%, 67.1%), emphasizing its critical role. Conversely, the reference points change penalty ($r_{t}^{c}$) slightly affects success rates but is significant for motion smoothness, where $r_{t2}$ shows increased Total Abs Acc (27600.56 in Stage 5) compared to $r_{t1}$(26475.46 in Stage 5), indicating smoother navigation with $r_{t}^{c}$ included. Overall, the results of the ablation study affirm the efficacy of our reward function design, highlighting the essential role of $r_{t}^{f}$ for attaining higher success rate and the contribution of $r_{t}^{c}$ to enhancing motion smoothness.

### 14 Mar 2024 v2

@@ III-E Observation Space Formulation
At time $t$, the observation for the agent $o_{t}$, as illustrated in Equation 5 and Figure 4, includes: the Euclidean distance $d_{t}$ from robot to navigation goal, the velocity magnitude $v_{t}$, the angular deviation $d\theta_{t}$ between robot’s velocity direction and the line to goal, the short-term and long-term reference points $Q_{t-1}^{s}$ and $Q_{t-1}^{l}$ from the previous frame’s action, the MPC-optimized points $Q_{1}^{*}$ and $Q_{N}^{*}$ (defined in Section III-D), and the convex region $convex_{t}$.
@@
Due to the variability in shape of the convex obstacle-free region ($convex_{t}$) derived from different point cloud data, the number of vertices in convex polygons is not constant. This poses a challenge for neural networks, which require fixed dimension input. To address this issue, the number of vertices in the calculated convex region, denoted as $num_{v}$ , is adjusted to match the predefined number $rnum_{v}$: vertices are added through interpolation if the count $num_{v}$ is less than $rnum_{v}$, or reduced through sparsification if $num_{v}$ exceeds $rnum_{v}$.
The agent can extract motion information about dynamic obstacles and its own historical trajectories from continuous frames implicitly. The agent’s state, denoted as $s_{t}$, a concatenation of observations from three consecutive frames $\left(o_{t-2},o_{t-1},o_{t}\right)$. The dimension of this state representation is set to $336rnumv+$ 33. This representation contains rich spatiotemporal information that is crucial for the agent’s decision making process.
@@ III-F Reward Function Formulation
The reward function plays a crucial role by setting the missions for agent. Agent’s policy is formulated based on the expectation of future rewards, making reward function a critical mechanism for conveying tasks to the agent.
A positive reward $r_{success}$ is awarded to the agent when its distance to the goal $d_{t}$ falls below $d_{th}$.
@@
The collision penalty $r_{t}^{o}$, derived from distance data $scan$ collected by LiDAR, is dynamically adjusted. The penalty decreases exponentially as the distance increases, controlled by a positive decay factor $w_{obs}$.
@@
To mitigate sparse reward, we introduce the reward $r_{t}^{a2}$ , based on the Euclidean distance $d_{t}$ between the robot and its destination.
@@
Short-term and long-term reference points, $Q_{t}^{s}$ and $Q_{t}^{l}$, discussed in Section III-C, might not be practical to reach. However, the MPC-optimized positions $Q_{i}^{*}$ , outlined in Section III-D, comply with kinematic constraints and provide more feasible references. Thus, to penalize the discrepancy between reference points and MPC-optimized points, we introduce $r_{t}^{f}$.This reward encourages the generation of reachable reference points.
@@
Significant changes in reference points across frames can result in deviations between MPC-optimized trajectories at consecutive steps. Consequently, the agent requires increased control efforts to mitigate the influence of previous frames. To address this, the penalty $r_{t}^{c}$ is introduced, which can also facilitate rapid learning of smooth direction changes during the initial training phase.
@@
The final reward function, denoted as $r_{t}$, is composed of the above six components. These components can be adjusted or selectively ignored during phases of training.
$r_{t}=r_{t}^{e}+r_{t}^{s}+r_{t}^{a2}+r_{t}^{o}+r_{t}^{c}+r_{t}^{f}$ (11)
@@ III-G Network Architecture
The convex obstacle-free region efficiently removes redundant information from the raw LiDAR point cloud, including noise, duplicate points, and distant obstacles. This preprocessing step allows for the implementation of a more streamlined network structure for fitting policy and value functions, as depicted in Figure 5. Given that separate networks can yield better performance in practice [27], the policy and value network are updated independently, without sharing parameters.
Figure 5: Network Architecture: The value network $V_{\phi}$ and policy network $\pi_{\theta}$ are structured as four-layer fully connected networks. The policy network is augmented with a logarithmic standard deviation parameter ($\ln\sigma_{t}$) for each action dimension. This configuration primarily facilitates the generation of stochastic policies by allowing actions to be sampled from a Gaussian distribution, where $\sigma_{t}$, derived from $\ln\sigma_{t}$, modulates the balance between exploration and exploitation.
@@ IV EXPERIMENTS
In this section, we introduce our training process along with the definition of seven Stages that differ in complexity (Table I). These stages are specifically designed for curriculum learning and subsequent evaluation. Following this, we evaluate our method in three parts:
(1) To demonstrate the advantages of our method in terms of action and state space design, we conduct an ablation study. This study compares our method with four different DRL navigation methods constructed with various action and state spaces (Section IV-B).
(2) To validate the effectiveness of our reward function design, we conduct another ablation study. This study conducts a comparison between our method and three other DRL navigation methods with different reward configurations (Section IV-C).
(3) In order to further analyze our method’s advantages over non-DRL navigation methods in dynamic and crowded environments, we conduct a comparison between our method and the TEB [28] in test scenarios (Section IV-D).
@@ IV-A Training Procedure
Our method employs a multi-stage training strategy that progresses from simple to complex scenarios, following the principle of curriculum learning. This strategy aims to enhance the agent’s learning efficiency by optimizing the sequence of acquiring experience. The agent first learns basic navigation movements in an obstacle-free environment (Stage 1) initially and gradually being exposed to increasingly complex static and dynamic obstacles.
@@ IV-A1 Stage Definition
The parameters of static and dynamic obstacles over seven Stages for staged training and subsequent evaluation are presented in Table I. Static obstacles are randomly generated polygons with three or four sides, where the maximum area of these polygons does not exceed 2 square meters. The dynamic obstacles are circular in shape.
In order to assess the performance of the trained agent, we generate 1,000 test scenarios for each Stage, from 2 to 7. These scenarios are generated by utilizing random seeds that are not included in the training dataset to configure the simulator.
@@ IV-A2 Implementation Details
Our study utilizes the Arena-Rosnav framework [29], developed within the ROS architecture, to enable reliable robotic navigation solutions. We also adopt the implementation of Proximal Policy Optimization (PPO) [30] from Elegant-RL [31], which provides tools for network customization and training process monitoring. The simulation experiments are conducted on a high-performance PC equipped with an Intel Core i7-7700K CPU, an Nvidia RTX 2080Ti GPU, and 32GB of RAM.
@@ IV-B Ablation Study on State and Action Space
To verify the effectiveness of the proposed method in terms of action and state space design, we conducted comparison experiments against four methods:
Design 1: Utilizing the conventional end-to-end architecture where the observation $o_{t}$ includes LiDAR data $scan_{t}^{{}^{\prime}}$ Action is defined as a set of two-dimensional velocities $\left(v_{x},v_{y}\right)$ The relationship between the output of the policy network $\left(\alpha_{t},\beta_{t}\right)$ and $\left(v_{x},v_{y}\right)$ is detailed in (12).
@@
Design 2: The state observation $o_{t}$ is updated to (13), the convex obstacle-free region is taken into account. The action, consisting of $(v_{x},v_{y})$, consistent with Design 1.
@@
Design 3: The action is simplified to $(Q_{t}^{l})$, focusing exclusively on the long-term reference point. The calculation process of a single $(Q_{t}^{l})$ is similar to the short-term reference point $(Q_{t}^{s})$ in Section III-C, except that single $(Q_{t}^{l})$ is sampled from the long-term action space. In addition, Design 3 simplifies the state observation in (14) by removing terms related to $Q_{t}^{s}$.
@@
Design 4: The observation $o_{t}$ includes raw LiDAR data $scan_{t}^{{}^{\prime}}$, as shown in (15). The action is still $\left(Q_{t}^{l},Q_{t}^{s}\right)$, but calculated within the intersection of the point cloud’s coverage and the robot’s kinematic limits. Given that $scan_{t}^{{}^{\prime}}$ is non-convex and cannot be used as a constraint for convex optimization, the constraint that the optimized trajectory must be within the convex region is removed during the MPC calculation.
@@
After training from Stage 2 to 4, the performance of our method and Designs 1 to 4 in test scenarios is shown in Table II.
@@
Stage Method Success Rate (%) Time (s) Distance (m) Speed (m/s)
Design 1 76 4.0 11.28 2.87
Design 2 76 4.0 11.11 2.8
2 Design 3 90.3 9.0 15.26 1.72
Design 4 83 5.0 11.66 2.23
Ours 89.2 5.5 12.19 2.18
Design 1 79.9 4.0 11.37 2.9
Design 2 79 4.0 11.4 2.88
3 Design 3 87.8 8.8 15.42 1.77
Design 4 83.5 4.9 11.59 2.24
Ours 89.1 5.9 13.04 2.17
Design 1 80.9 3.9 11.25 2.95
Design 2 75.1 3.8 10.78 2.89
4 Design 3 84.5 9.0 15.13 1.69
Design 4 83.7 4.8 11.19 2.2
Ours 85.5 5.8 12.51 2.1
In terms of action space design, Design 3 simplifies action to a single long-term reference point. Although Design 3 has an average navigation success rate of 87.53%, close to our method’s 87.93%, Design 3’s average navigation time is 3.2s longer and the average navigation distance is 2.69m longer than our method in the three Stages. On the other hand, compared with Design 1 and 2, the navigation success rate of our method is 9% and 11.23% higher respectively.
In terms of state space design, compared with Design 1 and Design 4 that directly use raw LiDAR data as the observation, our method’s average navigation success rate is 9% and 4.53% higher, respectively. The ablation study demonstrates the effectiveness of our method in the design of the action and state space.
@@ IV-C Ablation Study on Reward Function
The configuration of reward function $r_{t1}$ used in our method includes all items.
$r_{t1}=r_{t}^{e}+r_{t}^{s}+r_{t}^{a2}+r_{t}^{o}+r_{t}^{c}+r_{t}^{f}$ (16)
@@
Stage Reward Success Rate (%) Time (s) Distance (m) Speed (m/s) Total Abs Acc
5 $r_{t1}$ 87.1 2.9 4.59 1.57 26475.46
$r_{t2}$ 87.8 2.9 4.65 1.57 27600.56
$r_{t3}$ 84.8 3.4 4.38 1.40 27781.34
$r_{t4}$ 86.7 2.9 4.65 1.58 26460.66
6 $r_{t1}$ 78.1 2.9 4.12 1.42 19507.60
$r_{t2}$ 77.3 2.8 4.10 1.43 20397.66
$r_{t3}$ 73.1 3.4 3.85 1.25 21988.81
$r_{t4}$ 74.6 3.0 4.35 1.46 19950.77
7 $r_{t1}$ 68.5 3.1 4.22 1.36 6348.57
$r_{t2}$ 69.9 3.0 4.15 1.37 7218.53
$r_{t3}$ 67.1 3.6 3.98 1.19 6080.69
$r_{t4}$ 65.3 2.9 3.88 1.36 5668.30
When evaluating the impact of composite reward functions on motion smoothness, Total Absolute Acceleration (Total Abs Acc) is a crucial metric. Total Abs Acc is the sum of the agent’s absolute acceleration taken across steps within successful episodes.
The ablation study clarifies the importance of the MPC-optimized points error penalty ($r_{t}^{f}$) for high success rates. This is evident in the performance of $r_{t1}$’s performance, which achieved success rates of 87.1%, 78.1%, and 68.5% across stages. When $r_{t}^{f}$ is excluded in $r_{t3}$, the success rates decline to 84.8%, 73.1%, 67.1%. Conversely, the reference points change penalty ($r_{t}^{c}$) slightly affects success rates, but it is significant for motion smoothness. The $r_{t2}$ shows increased Total Abs Acc (27600.56 in Stage 5) compared to $r_{t1}$(26475.46 in Stage 5), indicating smoother navigation with $r_{t}^{c}$ included.